OpenStudy (anonymous):

Your school drama club is putting on a play and hopes to raise $600 from ticket sales. They sell adult tickets for $8 each and student tickets for $5 each. Let x represent the number of adult tickets and y represent the number of student tickets. The equation to represent the number of tickets sold to reach their goal, 8x+ 5y = 600, is graphed below. If 40 adult tickets are sold, how many student tickets must be sold to reach their goal?

OpenStudy (anonymous):

OK! Hold on. I'm solving :)

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

Ok. 56 kids.

OpenStudy (anonymous):

56 student tickets must be sold to reach the goal

OpenStudy (texaschic101):

8x + 5y = 600 8(40) + 5y = 600 320 + 5y = 600 5y = 600 - 320 5y = 280 x = 56 ANSWER : 56 student must be sold to reach their goal.

OpenStudy (anonymous):

Becauseee!!! Adult: $8 * 40 = 320. So you already have 320. 600 - 320 = 280. Divide that by 5, and you get 56 :) I was first to answer :)

OpenStudy (anonymous):

okay thank you! could you help me with this one also? please! The local gym charges members a monthly fee of $50 plus an additional $2 every time you want to use the pool. Let p represent the number of times you used the pool in one month and C represent the total cost for your monthly membership. Write an equation to find the total cost of your monthly membership to the gym.

OpenStudy (texaschic101):

she is right....she was first...lol

OpenStudy (anonymous):

I'm sorry! I gtg and can't help with the other one but I'm sure @texaschic101 could totally handle it :) :)

OpenStudy (texaschic101):

50 + 2p = C